OpenStudy (anonymous):

Difficult question, please help!! What is the simplified form of the quantity 3 x squared over x squared minus 1 all over 9 x over the x squared plus 2 x plus 1 ? a)x times the quantity x minus 1 all over 3 times the quantity x plus 1 b)x times the sum x plus 1 all over 3 times the quantity x minus 1 c)3 x times the quantity x plus 1 all over the quantity x minus 1 d)the quantity x minus 1 all over 3 x times the quantity x plus 1

OpenStudy (anonymous):

I tried to eliminate the choices, but I kept messing up. I just need help figuring out how to solve

OpenStudy (anonymous):

\[\frac{3x^2}{x^2-1}/\frac{9x}{x^2+2x+1} \]Is the above the problem expression?

OpenStudy (anonymous):

The first fraction that you put goes above the second.

OpenStudy (anonymous):

Oh I read that awrong. You're right!

OpenStudy (anonymous):

Well "/", although small, is a division sign.\[\left(\frac{3x^2}{x^2-1}/\frac{9x}{x^2+2x+1}\right)=\left(\frac{3x^2}{x^2-1}*\frac{x^2+2x+1}{9x}\right) \]Try factoring the expressions containing x^2. Also, 3x^2/9x will simplify.

OpenStudy (anonymous):